\begin{thebibliography}{WWW}
\bibitem{ako} M. Aschbacher, R. Kessar and B. Oliver, \textit{Fusion Systems in Algebra and Topology, London Math. Soc. Lect. Note Series}

\bibitem{caen} M. Cabanes and M. Enguehard. \textit{Representation Theory of Finite Reductive Groups, New Math. Monogr.} {\textbf 1}, Cambridge Univ. Press, Cambridge, 2004.

\bibitem{david} D. Craven, \textit{The Theory of Fusion Systems$:$ An algebraic approach, Cambridge Studies in Advanced Mathematics} {\textbf{131}}, Cambridge Univ. Press, Cambridge, 2011.

\bibitem{torquay} P. Fong and B. Srinivasan, {\em The blocks of finite general linear and unitary groups}, Invent. Math. {\bf 69}  (1982), no. 1, 109--153.

\bibitem{grazianpearls} V. Grazian, {\em Fusion systems containing pearls}, J. Algebra. {\bf 510} (2018), 98--140.

\bibitem{grazian} V. Grazian, {\em Fusion systems on $p$-groups of sectional rank $3$}, J. Algebra. {\bf 537} (2019), 39--78.

\bibitem{solom} R. Kessar. \textit{The Solomon system $\mathcal{F}_{\mathrm{Sol}}(3)$ does not occur as fusion system of a $2$-block.} J. Algebra {\textbf{296}}  (2006), no. 2, 409--425.

\bibitem{radha} R. Kessar. \textit{Introduction to block theory} in: “Group Representation Theory" (eds. M. Geck, D. Testerman, J. Thévenaz), EPFL Press, Lausanne (2007), 47--77.

\bibitem{ks} R. Kessar and R. Stancu, \textit{A reduction theorem for fusion systems of blocks}. J. Algebra {\bf 319}  (2008), no. 2, 806--823.

\bibitem{S20} P. Serwene,  {\em Block-exoticity of a family of exotic fusion systems},  Journal of Group Theory {\bf 23.3} (2020), 415--433.

\bibitem{ps23} P. Serwene,  {\em Reduction theorems for generalised block fusion systems},  J. Algebra {\bf 614} (2023), 458--480.

\bibitem{S24} P. Serwene,  {\em Proving a conjecture for fusion systems on a class of groups},  J. Algebra {\bf 638} (2024), 617--624.

%\bibitem{AB79} J. Alperin and M. Broué, {\em Local Methods in Block Theory}, Ann. Math. {\bf 110.1} (1979), 143–157.

%\bibitem{caen} M. Cabanes and M. Enguehard. \textit{Representation Theory of Finite Reductive Groups, New Math. Monogr.} {\textbf 1}, Cambridge Univ. Press, Cambridge, 2004.

%\bibitem{david} D. Craven. \textit{The Theory of Fusion Systems$:$ An Algebraic Approach, Cambridge Studies in Advanced Mathematics} {\textbf{131}}, Cambridge Univ. Press, Cambridge, 2011.

%\bibitem{derm} D.I. Deriziotis and G.O. Michler. \textit{Character table and blocks of finite simple triality groups $^3D_4(q)$.} Trans. Amer. Math. Soc. {\textbf{303}}  (1987), no. 1, 39--70.

%\bibitem{drv} A. Díaz, A. Ruiz and A. Viruel, \textit{All $p$-local finite groups of rank two for odd prime $p$}. Trans. Am. Math. Soc. {\bf 359}  (2007), no. 4, 1725--1764.

%\bibitem{gls} D. Gorenstein, R. Lyons and R. Solomon, {\em The Classification of Finite Simple Groups, Number 3, Math. Surveys Monogr.} {\bf 40}, American Mathematical Society, Providence, RI, 1998.

%\bibitem{hiss} G. Hiss and J. Shamash. \textit{$3$-blocks and $3$-modular characters of $G_2(q)$.} J. Algebra {\textbf{131}}  (1990), no. 2, 371--387.

%\bibitem{malle} G. Malle. \textit{Die unipotenten Charaktere von $^2F_4(q^2)$.} Commun. Algebra {\textbf{18}}  (1990), no. 7, 2361--2381.

%\bibitem{parkersemeraro} C.W. Parker, J. Semeraro,  {\em Fusion systems on maximal class $3$-groups of rank two revisited},  Proc. Am. Math. Soc. {\bf 147} (2019), no. 9, 3773--3786.

%\bibitem{GLS3}  D. Gorenstein, R.Lyons, R. Solomon, {\em  The Classification of Finite Simple Groups. Number 3. Part I},  vol. 40  of {\em Mathematical Surveys and Monographs}, American Mathematical Society, Providence, RI (1998).

%\bibitem{H} J.E. Humphreys, {\em  The Steinberg representation.} Mathematical Surveys and Monographs {\bf 16.2} (1987): 247--263.

%\bibitem{KLLS} R. Kessar, M. Linckelmann, J. Lynd, J. Semeraro,  {\em Weight conjectures for fusion systems}, Advances in Mathematics {\bf 357} (2019), 106825.

%\bibitem{KS} R. Kessar, M. Schaps. {\em Crossover Morita equivalences for blocks of the covering groups of the symmetric and alternating groups}, J. Group Theory {\bf 9} (2006): 715--730.

%\bibitem{OV} B. Oliver, J. Ventura, {\em Saturated fusion systems over $2$ groups}, Trans. AMS. {\bf 361} (2009), 6661--6728.

%\bibitem{PS18} C.W. Parker, J. Semeraro,  {\em Fusion systems over a Sylow $p$-subgroup of $G_2(p)$},  Mathematisches Zeitschrift {\bf 289} (2018), 629--662.


%\bibitem{Feitbook} W. Feit, {\em The Representation Theory of Finite
%Groups}, North-Holland Mathematical Library {\bf 25}, North-Holland
%Publishing Company, Amsterdam (1982).


%\bibitem{Gor} D. Gorenstein, {\em Finite Groups},
%Chelsea Publishing Company, New York (1980).



%\bibitem{LiBookI} M. Linckelmann, {\em The block theory of finite group
%algebras I}, Cambridge University Press, London Math. Soc. Student
%Texts {\bf 91} (2018). 

%\bibitem{LiBookII} M. Linckelmann, {\em The block theory of finite group
%algebras II}, Cambridge University Press, London Math. Soc. Student
%Texts {\bf 92} (2018). 


\end{thebibliography}

\documentclass{amsart}
\usepackage[T1]{fontenc} 

\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{xcolor}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage[utf8]{inputenc}
\usepackage[numbers]{natbib}
\usepackage{graphicx}
\usepackage{mathtools}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{tikz-cd} 
\usepackage{enumerate}
\usepackage[english]{babel}
\usepackage{hyperref}
\usepackage{tikz}
\usetikzlibrary{positioning} 
\usepackage{nomencl}
\usepackage[matrix, arrow]{xy}
\xyoption{arrow}
\theoremstyle{plain}\newtheorem{Theorem}{Theorem}[section]
\theoremstyle{plain}\newtheorem{Conjecture}[Theorem]{Conjecture}
\theoremstyle{plain}
\theoremstyle{plain}
\theoremstyle{plain}\newtheorem{Proposition}[Theorem]{Proposition}
\theoremstyle{definition}
\theoremstyle{definition}
\theoremstyle{definition}
\theoremstyle{definition}
\theoremstyle{definition}
\theoremstyle{definition}
\theoremstyle{definition}

\def\sp{\mathrm{Sp}}

\def\Syl{\mathrm{Syl}}

\DeclareMathOperator{\su}{SU}

\DeclareMathOperator{\slg}{SL}

\DeclareMathOperator{\gu}{GU}

\DeclareMathOperator{\mor}{Mor}

\DeclareMathOperator{\syl}{Syl}
\DeclareMathOperator{\gl}{GL}

\makeindex
\title{Exotic Fusion System on a Subgroup of the Monster} 
\author{Patrick Serwene}
\address{Technische Universität Dresden, Faculty of Mathematics, 01062 Dresden, Germany}
\email{patrick.serwene@tu-dresden.de}

\date{\today}

\begin{document}

\begin{abstract}
We prove that an exotic fusion system described by Grazian on a subgroup of the Monster group is block-exotic, thus proving that exotic and block-exotic fusion systems are the same for all $p$-groups with sectional rank $3$, where $p \geq 5$. 
\end{abstract}

\maketitle
\section{Introduction}
Consider a prime number $p$ and a finite $p$-group $P$. We define a \textit{fusion system} as a category where the objects are the subgroups of $P$, referred to as a category on $P$, and the morphisms are injective group homomorphisms between these subgroups, subject to specific conditions. If these morphisms adhere to two additional axioms, we term the fusion system \textit{saturated}. For simplicity, we shall henceforth use the term "fusion system" to mean "saturated fusion system."\\
Every finite group $G$ gives rise to a fusion system $\mathcal F_P(G)$ on a Sylow $p$-subgroup $P$ of $G$, where the morphisms are defined as conjugation maps induced by a fixed element in $G$, where they are well-defined. A fusion system constructed in this manner is termed \textit{realizable}, while one that cannot be constructed as such is termed \textit{exotic}. Let $k$ be an algebraically closed field with characteristic $p$, and let $b$ be a block of $kG$. In this context, we can also define a fusion system on a defect group $P$ of $b$ by defining the morphisms as well-defined conjugation maps induced by an element in $G$. This fusion system is denoted by $\mathcal F_{(P,e_P)}(G,b)$, where $(P,e_P)$ is a maximal $b$-Brauer pair. Not every fusion system $\mathcal F$ can be realized in this way; if it can, we call $\mathcal F$ \textit{block-realizable}; otherwise, it is termed \textit{block-exotic}. The following fact follows from Brauer’s Third Main Theorem (see \cite[Theorem 3.6]{radha}): If $G$ is a finite group and $b$ is the principal $p$-block of $kG$, i.e., the block corresponding to the trivial character, with maximal $b$-Brauer pair $(P,e_P)$, then $P \in \syl_p(G)$ and $\mathcal{F}_{(P,e_P)}(G,b)=\mathcal{F}_P(G)$. Hence, any realizable fusion system is block-realizable. However, the converse remains an open problem and has been conjectured for some time, see \cite[Part IV,7.1]{ako} and \cite[9.4]{david}:

\begin{Conjecture}
\label{XX}
\textit{If $\mathcal F$ is an exotic fusion system, then $\mathcal F$ is block-exotic.}
\end{Conjecture}

This conjecture is hard to tackle, since so far it has been mainly proved for one exotic family of fusion systems at a time, see \cite{solom}, \cite{david}, \cite{ks}, \cite{S20} or \cite{S24}. Since block fusion systems "misbehave" with regards to normal subgroups, a descend to normal subgroups and thus a reduction is not easily done and requires more general structures, see \cite{ks} or \cite{ps23} for an explanation.\\
In this paper, we prove Conjecture \ref{XX} for all fusion systems on $p$-groups with sectional rank $3$, where $p \geq 5$. A group has sectional rank $r$ if $r$ is the minimal integer such that any of its subgroups has at most rank $r$. In \cite{grazian}, fusion systems on $p$-groups of sectional rank $3$ are studied and if we further assume that $p \geq 5$ and $O_p(\mathcal F)=1$, Grazian proves in Theorem C that the only options for such a fusion system are either a unique exotic system on a group of order $7^5$, or the fusion system of $\sp_4(p)$ on its Sylow $p$-subgroup. The group of order $7^5$ hosting the exotic fusion system is a maximal subgroup of a Sylow $7$-subgroup of the Monster group. Here we prove that the fusion system is block-exotic too, giving rise to our main theorem:

\begin{Theorem}
\label{main}
\textit{Let $p \geq 5$ be a prime, $P$ a $p$-group of sectional rank $3$ and $\mathcal F$ be a fusion system on $P$ with $\mathcal O_p(\mathcal F)=1$. Then $\mathcal F$ is exotic if and only if it is block-exotic}.
\end{Theorem}

See \cite{ako} for details on (block) fusion systems or Section 2 of \cite{S20} for a more compact overview of the terms needed. In the only remaining section we prove our main theorem, using the Classification of Finite Simple Groups.

\section{Proof of Main Theorem}
Fix $P$ to be a maximal subgroup of a Sylow $7$-subgroup of the monster. We prove that the exotic fusion system on $P$ described by Grazian in \cite{grazian} is block-exotic too. We first reduce the problem to quasisimple groups and state the reduction theorem we apply. Recall that a fusion system is called reduction simple, when it has no non-trivial proper strongly $\mathcal F$-closed subgroups, i.e. subgroups $Q$ such that $\varphi(R) \leq Q$ for all $\varphi \in \mor(\mathcal F)$ defined on $R \leq Q$.

\begin{Theorem}
\cite[Theorem 3.5]{S20}
\label{glastonbury}
\textit{Let $P$ be a non-abelian $p$-group such that $Z(P)$ is cyclic and let $\mathcal F$ be a reduction simple fusion system on $P$. If $\mathcal F$ is block-realizable, then there exists a fusion system $\mathcal F_0$ on $P$ and a quasisimple group $L$ with an $\mathcal F_0$-block, where $\mathcal O_p(\mathcal F_0)=1$}.
\end{Theorem}

\begin{Proposition}
\label{redu}
\textit{Let $\mathcal F$ be the exotic fusion system on $P$. If $\mathcal F$ is block-realizable, it is block-realizable by the block of a finite quasisimple group}.
\end{Proposition}

\begin{proof} Clearly $P$ has cyclic centre. Assume $1 \neq N \leq P$ is strongly $\mathcal F$-closed. In particular $N \unlhd P$, which implies $Z(P) \leq N$. Thus, as in the proof of \cite[Theorem 4.3.1]{grazianpearls}, we obtain $N=P$, which means that $\mathcal F$ is reduction simple. By Theorem \ref{glastonbury}, there exists a fusion system $\mathcal F_0$ on $P$ and a quasisimple group $L$ with an $\mathcal F_0$-block, where $\mathcal O_p(\mathcal F_0)=1$. However, by Theorem C of \cite{grazian} we deduce $\mathcal F=\mathcal F_0$.
\end{proof}

The following result is useful for proving block-exoticity.

\begin{Proposition}
\label{bologna}
\cite[Proposition 4.3]{S20}
\textit{Let $G$ be a quasisimple finite group and denote the quotient $G/Z(G)$ by $\overline G$. Suppose $\overline{G}=G(q)$ is a finite group of Lie type and let $p$ be a prime number $\geq 7$, $(p,q)=1$. Let $D$ be a $p$-group such that $Z(D)$ is cyclic of order $p$ and $Z(D) \subseteq [D,D]$. If $D$ is a defect group of a block of $G$, then there are $n, k \in \mathbb{N}$ and a finite group $H$ with $\slg_n(q^k) \leq H \leq \gl_n(q^k)$ $($or $\su_n(q^k) \leq H \leq \gu_n(q^k))$ such that there is a block $c$ of $H$ with non-abelian defect group $D'$ such that $D'/Z$ is of order $|D/Z(D)|$ for some $Z \leq D' \cap Z(H)$}.
\end{Proposition}

\begin{Proposition}
\label{hjelp}
\textit{If $G$ is as in the previous proposition, then $G$ has no blocks with defect groups isomorphic to $P$}.
\end{Proposition}

\begin{proof} Recall $|P|=7^5$ and $Z(P) \cong C_7$. We apply Proposition \ref{bologna} with $P$ taking the role of $D$. Let $H$, $D'$ be as in its assertion with $p=7$. Assume first $H \leq \gl_n(q^k)$ and let $a$ be such that $|q^k-1|_7=7^a$. Then, since $\slg_n(q^k) \leq H$, we have $|D'|=|P/Z(P)| \cdot |Z|=7^4|Z| \leq 7^4|Z(H)| \leq 7^4|Z(\slg_n(q^k))|\leq 7^{4+a}$. Now the block of $k\gl_n(q^k)$ covering $c$ has a defect group of order at most $7^{2a+4}$. But it is a well-known fact, that (non-abelian) defect groups of $\gl_n(q^k)$ have order at least $7^{7a+1}$, see \cite[Theorem 3C]{torquay}. Thus, $7^{7a+1} \leq 7^{2a+4}$, which is a contradiction. The case $H \leq \gu_n(q^k)$ can be shown in the same fashion by considering the $7$-part of $q^k+1$ instead of $q^k-1$. \end{proof}

\textit{Proof of Theorem \ref{main}}.
By \cite[Theorem C]{grazian}, the only exotic fusion system on such a group is a fusion system on a maximal subgroup $P$ of a Sylow $7$-subgroup of the Monster with $|P|=7^5$. Thus, by \cite[Theorem 3.6]{radha}, we only need to prove block-exoticity of this system, denote it by $\mathcal F$.\\
By Proposition \ref{redu}, if $\mathcal F$ is block-realizable, we may assume that it is block-realizable by the block of a finite quasisimple group $G$ having $P$ as defect group. We use the classification of finite simple groups to exclude all possibilites for $G$. Clearly, we can assume that $G$ is non-abelian.\\
Firstly, assume $G/Z(G)$ is an alternating group $\mathfrak A_m$. Then $P$ is isomorphic to a Sylow $7$-subgroup of some symmetric group $\mathfrak S_{7w}$ with $w \leq 6$. Define the cycle $\sigma_i=((i-1)7+1,(i-1)7+2,\dots,i7)$ and the subgroup $S'=\langle \sigma_1, \dots, \sigma_6 \rangle\leq \mathfrak A_m$. Then $S' \in \syl_7(\mathfrak A_m)$. But this group is abelian, which means that $P \notin \Syl_7(\mathfrak A_m)$ and thus it cannot be the defect group of a block of $G$.\\
Next, assume $G$ is a group of Lie type. First assume the latter group is defined over a field of characteristic $7$, then $P$ is a Sylow $7$-subgroup of $G$ by \cite[Theorem 6.18]{caen} and thus $\mathcal F$ cannot be exotic. In particular, we can assume $G$ is defined over a field of order coprime to $7$. However, by Proposition \ref{hjelp}, we see that case in cross characteristic is not possible either.\\
Finally, assume $G/Z(G)$ is one of the sporadic groups. By, \cite[Theorem 9.22]{david}, the fusion system of a block of a sporadic group can not be exotic. This proves the theorem. \hfill$\square$\\

\bigskip

{\it Acknowledgements.} I thank Chris Parker and the University of Halle--Wittenberg to provide a platform to meet him. Furthermore, I thank the referee for carefully reading my manuscript and Pham Huu Tiep for communicating this paper.


\end{document}